### v2

# Integrating Random Regret Minimization-Based Discrete Choice Models with Mixed Integer Linear Programming for Revenue Optimization

Abstract
This paper explores the critical domain of Revenue Management (RM) within Operations Research (OR), focusing on intricate pricing dynamics. Utilizing Mixed Integer Linear Programming (MILP) models, the study enhances revenue optimization by considering product prices as decision variables and emphasizing the interplay between demand and supply. Recent advancements in Discrete Choice Models (DCMs), particularly those rooted in Random Regret Minimization (RRM) theory, are investigated as potent alternatives to established Random Utility Maximization (RUM) based DCMs. Despite the widespread use of DCMs in RM, a significant gap exists between cutting-edge RRM-based models and their practical integration into RM strategies. The study addresses this gap by incorporating an advanced RRM-based DCM into MILP models, addressing pricing challenges in both capacitated and uncapacitated supply scenarios. The developed models demonstrate feasibility and offer diverse interpretations of consumer choice behavior, drawing inspiration from established RUM-based frameworks. This research contributes to bridging the existing gap in the application of advanced DCMs within practical RM implementations.
keywords:
Operations Research, Pricing, Demand Prediction, Mixed Integer Linear Programming, Choice-based Optimization, Decision-making Behavior
[…]
organization=Department of Integrated Systems Engineering, addressline=The Ohio State University, city=Columbus, postcode=43210, state=OH, country=USA
[…]
organization=School of Computing, addressline=Clemson University, city=Clemson, postcode=29632, state=SC, country=USA
[…]
organization=ENSIMAG, addressline=Grenoble Alpes University, city=Grenoble, postcode=38400, state=Saint-Martin d’Hères, country=France
1 Introduction
Advancements in Operations Research (OR) Modeling, coupled with innovative solving methods, have efficiently addressed large-scale challenges across economics, transportation, supply chain, revenue management, etc. The primary goal of these models is to optimize revenue, minimize costs, or mitigate pollution. Consequently, sophisticated demand models are under development to precisely quantify these factors. Recognizing the intricate interplay between demand and supply, our approach integrates both aspects, emphasizing simultaneous consideration of the supply side during supply chain planning. Notably, the pricing decision involves suppliers setting prices while buyers determine their willingness to pay, exemplified by the varying affordability of a \$10 rose in different neighborhoods. This underscores the crucial need for sellers to comprehend and anticipate customer purchasing behavior. Concurrently, Revenue Management (RM), within the realm of OR, concentrates on maximizing revenue through strategic pricing and capacity allocation. Our work uniquely combines economic and OR perspectives, incorporating customer behavior modeling for optimized supply systems.
Historically, demand and revenue-maximizing prices were determined using simple linear demand functions. However, this approach lacks accuracy due to individual variations in purchase preferences and willingness to pay. RM addresses these challenges by predicting customer behavior to enhance revenue talluri2006theory .
An essential assumption in RM is the variability of product prices for each customer. Precise understanding of customer behavior is necessary to optimize revenue. Discrete Choice Models (DCMs), which elucidate and predict customer choices, significantly contribute to customer behavior modeling. These models consider customer preferences, leading to efficient resource utilization and substantial revenue improvements. Numerous empirical studies underscore the impact of discrete choice modeling in diverse sectors such as RM talluri2004revenue , labor market kornstad2007discrete , food industry gracia2008demand , and airline industry vulcano2010om .
Random Utility Maximization Theory (RUM), established by mcfadden1978modelling ; mcfadden2000mixed , forms the foundation for discrete choice modeling. The core assumption of RUM is that decision-makers are perfectly rational and aim to maximize their utilities. Recent studies, such as those by chorus2008random ; chorus2010new ; chorus2012rrm propose an alternative approach by investigating DCMs derived from Random Regret Minimization (RRM) where the customer tries to minimize the regret caused by not choosing an alternative. Conditions exist where DCMs derived from either theory outperform the other, revealing diverse customer preferences hensher2013random ; greene2011ru ; li2018hybrid .
While existing literature often examines demand and supply separately or investigates supply based on a simplistic demand formulation, a comprehensive approach is needed for effective revenue generation pacheco2016new . Neglecting the interplay between supply and demand, or focusing solely on one side, can lead to suboptimal pricing strategies, resulting in reduced gains, the bullwhip effect, and discordance in supply chains. Recognizing that customer decisions drive demand, it underscores the vital importance of optimizing and coordinating both supply and demand concurrently. In this regard, pacheco2017integrating innovatively crafted a framework that characterizes how demand responds to suppliers’ decisions, linking supply and demand through price as a crucial variable. They linearized DCMs into predictors of demand using a Mixed Integer Linear Programming (MILP) model, emphasizing price variables in the objective function.
In related literature, other works have incorporated DCMs in OR to forecast disaggregated demand and maximize revenue. For example, pacheco2016new ; paneque2017integrating addressed challenges arising from the non-linearity of DCM models in optimization. To address this, utility functions (RUM-based) were integrated into MILP constraints, introducing individual utility functions as scores to be maximized in discrete optimization, crucial for characterizing the demand side of the OR problem.
Existing literature in RM predominantly estimates demand based on RUM theory. However, RRM theory, a recent challenger to RUM, has garnered significant attention among researchers. To our knowledge, there is currently no choice-based optimization method grounded in RRM. Creating a framework akin to pacheco2017integrating but rooted in RRM would empower researchers to evaluate both demand and supply concurrently, offering a distinct interpretation of customer behavior.
In this paper, a novel MILP model based on RRM and the work of pacheco2016new ; pacheco2017integrating has been introduced in the realm of choice-based optimization. This framework serves to connect economic theories with the optimization of supply chains.
The paper is structured as follows:
In Section 2, a comprehensive literature review is provided, encompassing revenue management, microeconomic behavioral theories, discrete choice modeling, choice-based optimization framework, and data collection. Section 3 introduces a novel representation of demand and supply based on RRM through MILP formulation. Section 4 conducts experiments, and Section 5 concludes.
2 Related Work
Within RM, pricing decisions hold prominent importance among various aspects to increase the revenue. Several RM techniques aim to leverage customer heterogeneity for a product set at a specific time. They often focus on fixing alternatives or products and time dimensions while exploring customer differences. In our context, we emphasize the significance of customer heterogeneity in purchasing behavior. If all customers perceive a product uniformly with identical purchasing behavior, the customer dimension diminishes, leading to unexploited variations in willingness to pay. These variations, influenced by socio-economic factors like marital status, salary, age, gender, etc., play a critical role in maximizing revenue. On the other hand, optimizing revenue hinges on accurate demand prediction, crucial for effective supply planning and pricing strategies. The demand function is influenced by various variables, encompassing customer heterogeneity, prices, etc talluri2006theory ; SADRALASHRAFI201882 . Specifically, researchers have extensively focused on DCMs, particularly due to their capacity to predict demand by accounting for consumer heterogeneity in tastes and attributes talluri2006theory ; mcfadden2000mixed ; chorus2008random . RUM-based DCMs assume decision-makers are rational utility maximizers, and models derived from these theories aim to reveal influential factors shaping customer choice preferences (e.g., see kemperman2021review for a recent review). However, in some cases, customers prefer regret minimization over utility maximization when making choices. Therefore, deriving DCMs from suitable behavioral theories is crucial hensher2013random . In this regard, chorus2014random introduced a regret-based discrete choice model for consumer preference analysis with multinomial choice sets. Their empirical assessment, based on 43 comparisons between RUM and RRM models in the literature, revealed the superiority of RRM-based DCMs over RUM-based ones. The study demonstrated that both RRM-based and hybrid RRM-RUM-based DCMs consistently outperformed their RUM-based counterparts in terms of prediction accuracy and overall model fit. The authors emphasized that although performance differences were minor, the models led to significant managerial implications and diverse interpretations of behavior.
The reference masiero2019understanding conducted a choice experiment to investigate tourists’ hotel preferences. They created hypothetical hotel alternatives and randomized factors, including price, hotel rating, and walking time to key attractions, to form choice sets. RRM and RUM-based DCMs were utilized to estimate the coefficients of these factors where they figured out the RRM-based DCMs performed significantly better than their RUM-based counterparts. Additionally, they showed that quantitative simulations can be performed to study the demand and pricing issues while incorporating customers’ heterogeneity into the model.
The reference andersson1998passenger made a pioneering contribution to Revenue Management by focusing on modeling passenger preferences using DCMs. Following andersson1998passenger , gallego2004managing addressed a general question in RM related to consumer choice behavior (RUM). They explored the network revenue management problem with flexible products, considering scenarios where demand was either exogenously given or generated by a consumer choice model based on RUM-based DCM. Their findings indicated that the optimal value of the stochastic optimization problem closely approximated that of a deterministic linear program in both settings.
The reference liu2008choice delved into the implications of the concept introduced by gallego2004managing in RM. They developed a Choice-based Deterministic Linear Programming (CDLP) model under consumer choice behavior, focusing on providing suitable offer sets, seats, and flights in the airline industry while adhering to capacity constraints over time. Their study demonstrated that the CDLP, inspired by gallego2004managing , serves as a deterministic approximation of the stochastic formulation, with its solution being an asymptotic optimal solution for the original stochastic problem influenced by the stochasticity of customer choice probability.
The reference bront2009column expanded on the CDLP introduced by liu2008choice . They addressed the issue of overlapping offer sets in the airline industry, where airlines typically present affordable morning tickets in one set and pricier afternoon tickets in another set. This extension considered combinations such as expensive and morning tickets, providing insights into more realistic scenarios. They applied column generation method to tackle the dimensionaliy issue of the problem. Particularly, meta-heauristc methods seem to be a promising approach in approximating high quality solutions for large-sized problems (e.g., see Haeri_Boroujeni_2023 ; mehrabi2023efficient ; boroujeni2021data ; 9721496 ; https://doi.org/10.13140/rg.2.2.29416.03849 ).
The reference azadeh2015impact introduced a modified CDLP model, solving it on a synthetic dataset to obtain optimal offer sets and an upper bound on revenue. Additionally, they suggested a non-parametric model for estimating choice probabilities, utilizing the synthetic data employed in the CDLP. References bierlaire2016demand ; PACHECOPANEQUE202126 ; PACHECOPANEQUE2022105985 introduced the linearized formulation of demand based on customer choice behavior, employing a multinomial logit model, a type of DCM which assumes the unobserved utility of the customer is independent of the attributes of the customer, and acknowledging the non-linearity and non-convexity of the resulting model. They integrated it into a MILP model, marking the emergence of RUM-based MILP models in RM. Subsequently, pacheco2017integrating linearized a mixed multinomial logit DCM model, a model where the unobserved utility of the customer is deemed to be correlated with the attributes of the customer, through MILP, aiming to maximize revenue by simultaneously exploring supply (price and capacity) and demand (price, utility levels, customer satisfaction) factors.
The reference paneque2019passenger delved into an economics-oriented approach, emphasizing passenger satisfaction maximization by integrating DCMs into MILP models. The objective was to enhance the public transportation network by utilizing toll revenue from highways, ensuring a balance between economic considerations and passenger contentment.
Generally, machine algorithms can be broadly categorized into one of the following three types: supervised, unsupervised, and agent-based learning. Paper boroujeni2024comprehensive provides a comprehensive overview of the existing ML techniques along with their types, characteristics and potential applications. Indeed, machine learning (ML) tools such as MNL and MMNL are used to predict customers’ choices based on RUM theory. For a recent literature on this regard, see hillel2021systematic ; soleymani_forecasting_2024 ; boroujeni2024ic ; 9814765 ; yousefpour2023unsupervised . Reinforcement learning as an online learning method can be also leveraged to help RM. For a recent literature on this regard, see jebellat2023reinforcement ; ran2024development ; amiri2023rank ; amirimargavi2023lowrank .
In conclusion, DCMs are state-of-the-art for estimating dis-aggregate demand, showing promise in RM. The literature demonstrates integrating DCMs into optimization models for better supply-demand interactions. However, a gap exists between DCMs derived from theories like RRM and optimization formulations. Finally, pacheco2016new ; pacheco2017integrating ; PACHECOPANEQUE202126 ; PACHECOPANEQUE2022105985 developed a RUM-based DCM framework where customers are utility maximizers. We extended the framework for RRM-based DCM introduced by chorus2010new , where the customers are regret minimizers.
3 Modelling
3.1 Problem definition
Random regret minimization asserts that individuals choose alternatives to minimize the regret they may experience from foregone options. Randomness arises from the inability of researchers to observe all the regret a customer perceives. To illustrate the RRM mechanism, consider the example of passenger travel mode choice provided by chorus2008random . The alternatives are car (1), train (2), and bus (3), each with attributes $x_{1},\ x_{2}$ and $x_{3}$ representing time, cost, and alternative-specific dummy variables of travel. The regret for each alternative is calculated, and a rational customer would choose the alternative with the minimum regret:
$R_{i}=\sum_{j\neq i}R_{ij}\ \text{and}\ R_{ij}=\sum_{k\in\{x_{1},x_{2},x_{3}\}}R_{ijk},\ \text{and}\ R_{ijk}=\max\{v_{ok},\beta_{k}(k_{j}-k_{i})+v_{k}\}$
Where $\beta$ parameters, representing consumers’ tastes for each attribute $k$, are estimated from historical data, and $k_{i}$ represents the $k^{th}$ attribute of alternative $i$. Regret, partially known to researchers, is modified by adding an error term, drawn from an independent and identically distributed (iid) extreme value type I distribution, to obtain the overall regret and $v_{ok}$ and $v_{k}$ stand for unobserved regret in attribute perceptions chorus2010new . Additionally, let $I=|C|$ represent the number of alternatives in the universal choice set $C$, indexed by $i$, where $i=0$ denotes the opt-out option when the customer chooses not to buy any alternatives. The set $N$ represents customers, indexed by $n$. For each customer, the choice set $C_{n}$ is considered, with each product $i$ available in quantity $c_{i}$. The supplier aims to maximize revenue by making decisions on whether a product $i$ should be offered to customer $n$ and what price of alternative $i$ to be offered to customer $n$, denoted as $p_{in}$. Customers are served sequentially based on a priority list, determined by factors such as customer loyalty, arrival order, customer segments, anticipated order quantities, etc. Such an assumption is standard (e.g., see pacheco2016new ; PACHECOPANEQUE202126 ). As it was shown by pacheco2016new ; paneque2017integrating , when solving this problem directly using DCMs, highly non-linear and non-convex problems arise. Therefore, to address this challenge, we employ MILP models. This paper explores two cases, the first assumes the supplier has no capacity limitation on supplies, while the second considers a scenario where there is a capacity limitation on the supply.
3.2 Model formulation
We have summarized the notations used in the model in the following.
3.2.1 Sets
$I=|C|$ Number of alternatives, indexed by $i$ ($i=0$ denotes the no-purchase
alternative or opt-out option)
$L_{in}$ Number of binary variables characterizing the price levels for each alternative
and individual, indexed by $I$
$N$ Number of customers, indexed by $n$
$R$ Number of draws, indexed by $r$
3.2.2 Input Parameters
$c_{i}$ Available quantity of alternative $\text{Available quantity of alternative }i\text{ for sale, maximum capacity of service }i$
[…]
$pm$ Price multiplier, to get desired price levels. Also, to set the price levels and
ranges precisely, one can increase or decrease the value of the multiplier
$ER_{ijnr}$ Sum of attribute level regret associated with alternative $i$ when compared
with alternative $j$ for customer $n$ and scenario $r$. Note that:
$ER_{ijnr}=\sum_{m=1}max\{v_{onmr},\beta_{nm}(x_{jm}-x_{im})+v_{xnmr}\}$
where $v_{onmr},v_{xnmr}$ are draws of Gumbel$(0,1)$. $x_{jm}$ represents the $m^{th}$ level of
alternative $j$’s attribute $x$. It is worth mentioning that price related regret has been
excluded from $ER_{ijnr}$
$mm_{ijnr}$ $=\max\{v_{onr},\beta_{np}(lp_{jn}-mp_{in})+v_{nr}\}\ \text{are upper bounds on variables $RR_{ijnr}$}$
$ll_{ijnr}$ $=\min\{v_{onr},\beta_{np}(mp_{jn}-lp_{in})+v_{nr}\}\ \text{are lower bound on variables $RR_{ijnr}$}$
[…]
$m_{inr}$ $=\sum_{j\neq i}\max\{v_{onr},\beta_{np}(lp_{jn}-mp_{in})+v_{nr}\}+ER_{ijnr},\ \text{upper bounds on variables}\ R_{inr}$
$l_{inr}$ $=\sum_{j\neq i}\min\{v_{onr},\beta_{np}(mp_{jn}-lp_{in})+v_{nr}\}+ER_{ijnr},\ \text{lower bounds on variables}\ R_{inr}$
$m_{nr}$ $=\max_{i\in C_{n}}\{m_{inr}\}\ \text{upper bounds on variables }R_{nr}$
$l_{nr}$ $=\min_{i\in C_{n}}\{l_{inr}\}\ \text{lower bounds on variables }R_{nr}$
[…]
$\beta_{nm}$ Customer $\text{Customer }n\text{ taste for attribute }m\text{ of alternatives}$
3.2.3 Variables
$y_{in}\in\{0,1\}$ Supplier’s decision when alternative $i$ is in the customer $n$’s choice set
$y_{inr}\in\{0,1\}$ Availability variables, $y_{inr}=1$ if alternative $i$ is available to customer $n$ at scenario $r$
$RR_{ijnr}$ Regret of customer $n$ for the price attribute of alternative $i$ when compared
with that of alternative $j$
$R_{inr}$ Overall regret of alternative $\text{Overall regret of alternative }i$
$R_{nr}$ Minimum regret among all available alternatives, i.e., $\min_{i\in C_{n}}R_{inr}$
$b_{ijnr}\in\{0,1\}$ Binary variables to capture the maximum of $RR_{ijnr}$
$z_{ijnr}$ Discounted regret, $z_{ijnr}\geq 0$ and $z_{ijnr}=RR_{ijnr}+ER_{ijnr}$ when $y_{jnr}=1$ and 0 otherwise
[…]
3.2.4 Uncapacitated supply MILP model
$\max\frac{1}{R}\sum_{n}\sum_{i>0,i\in C_{n}}\sum_{r}(lp_{in}w_{inr}+\sum_{l}l\alpha_{inrl}pm)$ (1)
$s.t.$
$v_{onr}\leq RR_{ijnr},$ $\forall i,j\in C_{n},i\neq j,n,r,$ (2)
$RR_{ijnr}\leq v_{onr}+M_{ijnr}b_{ijnr},$ $\forall i,j\in C_{n},i\neq j,n,r,$ (3)
$\beta_{np}(lp_{jn}+\sum_{l}pm\lambda_{jnl}l-lp_{in}-\sum_{l}pm\lambda_{inl}l)$
$+v_{nr}\leq RR_{ijnr},$ $\forall i,j\in C_{n},i\neq j,n,r,$ (4)
$RR_{ijnr}\leq\beta_{np}(lp_{jn}+\sum_{l}l\lambda_{jnl}pm-lp_{in}-\sum_{l}l\lambda_{inl}pm)$
$+v_{nr}+M_{ijnr}(1-b_{ijnr}),$ $\forall i,j\in C_{n},i\neq j,n,r,$ (5)
$R_{inr}=\sum_{j\neq i,j\in C_{n}}RR_{ijnr}+ER_{ijnr},$ $\forall i\in C_{n},n,r,$ (6)
$R_{nr}\leq R_{inr},$ $\forall i\in C_{n},n,r$ (7)
$R_{inr}-M_{nr}(1-w_{inr})\leq R_{nr},$ $\forall i\in C_{n},n,r$ (8)
$\alpha_{inrl}\leq\lambda_{inl},$ $\forall i\in C_{n},i\neq 0,n,r,l,$ (9)
$\alpha_{inrl}\leq w_{inr},$ $\forall i\in C_{n},i\neq 0,n,r,l,$ (10)
$\sum_{l}\lambda_{inl}\leq 1,$ $\forall i\in C_{n},i\neq 0,n,$ (11)
$lp_{in}+\sum_{l}l\lambda_{inl}pm\leq pm_{in}$ $\forall i\in C_{n},i\neq 0,n$ (12)
There are $R$ scenarios for each customer in the objective function 1, and the average revenue will be computed by dividing the total revenue over $R$, and $lp_{in}$ is a lower bound on the price levels. For an alternative $i$ and customer $n$, exactly one price level should be offered in scenario $r$. If alternative $i$ is chosen ($w_{inr}=1$), the price level should be selected ($\lambda_{inl}=1$) in a way that $lp_{in}+\sum_{l}l\alpha_{inrl}pm$ has the most positive contribution to the objective function while respecting constraints of the model. We define $pm$ as a constant price multiplier, and integers $l$ to characterize price levels. The opt-out alternative is excluded from the objective function as it does not affect the revenue. Constraints 2,3, 4, and 5 are to calculate the anticipated regret arising from the comparison of the price attribute of two alternatives. In this model, all alternatives are available for each customer. An error term $v_{nr}$ has been added to the price-raised regret between alternatives $i$ and $j$, $RR_{ijnr}$, to unveil the maximum anticipated regret.
The parameters $B_{np}$ shows customer $n$ taste for the price attribute, assumed to be negative for all customers. However, considering the RUM-based MILP, we do not compare alternatives pairwise to calculate regret. Instead, we measure the regret obtained from different price levels for each alternative. Then, the alternative rendering the minimum regret will be chosen. This is the major difference between the RUM-based and RRM-based MILPs.
The terms $ER_{ijnr}$ are the sum of the anticipated regret of attributes (other than the price attribute) for alternatives, and constraints 6 are to obtain the overall anticipated regret of an alternative using these terms and price attribute regret.
Constraints 7 and 8 determine the alternative with the lowest regret among all other alternatives in the choice set of a customer at each scenario. Here, the alternative providing the least regret will be chosen, while in RUM-based models, the one with the highest utility is chosen.
Constraints 9, 10 and 11 have been used to linearize the objective function. Constraints 12 illustrates the price bounds for each single customer at each scenario. $mp_{in}$ stands for upper bounds on the price, and $lp_{0n}=mp_{0n}=0$.
The rest of the constraints are bound constraints where the bounds were defined in the notation part:
$RR_{ijnr}\leq mm_{ijnr}$ $\forall i,j\in C_{n},i\neq j,n,r,$ (13)
$ll_{ijnr}<RR_{ijnr}$ $\forall i,j\in C_{n},i\neq j,n,r,$ (14)
$R_{inr}\leq m_{inr}$ $\forall i,j\in C_{n},n,r,$ (15)
$l_{inr}<R_{inr}$ $\forall i\in C_{n},n,r,$ (16)
$R_{inr},RR_{ijnr}\in\mathbb{R}$ $\forall i,j\in C_{n},i\neq j,n,r.$ (17)
These constraints typically enhance the efficiency of the model-solving process.
3.2.5 Capacitated supply MILP model
A capacitated version of the previously presented model has been depicted in this part. The term ”capacitated” is used to emphasize that both resources and the available quantity of an alternative are limited.
$\max\frac{1}{R}\sum_{n}\sum_{i\neq 0,i\in C_{n}}\sum_{r}(lp_{in}w_{inr}+\sum_{l}l\alpha_{inrl}pm)$ (18)
$s.t.$
$y_{in}=0,$ $\forall i\notin C_{n},n,r,$ (19)
$y_{inr}\leq y_{in},$ $\forall i,n,r,$ (20)
$v_{onr}\leq RR_{ijnr},$ $\forall i,j\in C_{n},i\neq j,n,r,$ (21)
$RR_{ijnr}\leq v_{onr}+M_{ijnr}b_{ijnr},$ $\forall i,j\in C_{n},i\neq j,n,r,$ (22)
$\beta_{np}(lp_{jn}+\sum_{l}l\lambda_{jnl}pm-lp_{in}-\sum_{l}l\lambda_{inl}pm)$
$+v_{nr}\leq RR_{ijnr},$ $\forall i,j\in C_{n},i\neq j,n,r,$ (23)
$RR_{ijnr}\leq\beta_{np}(lp_{jn}+\sum_{l}l\lambda_{jnl}pm-lp_{in}$
$-\sum_{l}l\lambda_{inl}pm)+v_{nr}+M_{ijnr}(1-b_{ijnr}),$ $\forall i,j\in C_{n},i\neq j,n,r,$ (24)
$z_{ijnr}\leq RR_{ijnr}+ER_{ijnr}$ $\forall i,j\in C_{n},i\neq j,n,r,$ (25)
$RR_{ijnr}+ER_{ijnr}-M(1-y_{jnr})\leq z_{ijnr}$ $\forall i,j\in C_{n},i\neq j,n,r,$ (26)
$z_{ijnr}\leq My_{jnr}$ $\forall i,j\in C_{n},i\neq j,n,r,$ (27)
$-My_{jnr}\leq z_{ijnr}$ $\forall i,j\in C_{n},i\neq j,n,r,$ (28)
$R_{inr}=\sum_{j\neq i,j\in C_{n}}z_{ijnr},$ $\forall i\in C_{n},n,r,$ (29)
$w_{inr}\leq y_{inr},$ $\forall i\in C_{n},n,r,$ (30)
$R_{nr}\leq R_{inr}+M_{nr}(1-y_{inr}),$ $\forall i\in C_{n},n,r$ (31)
$R_{inr}-M_{nr}(1-w_{inr})\leq R_{nr},$ $\forall i\in C_{n},n,r$ (32)
$y_{i(n+1)r}\leq y_{inr},$ $\forall i\in C_{n},i\neq 0,n<N,r,$ (33)
$\sum_{m=1}^{n-1}w_{imr}\leq(c_{i}-1)y_{inr}+(n-1)(1-y_{inr}),$ $\forall i\in C_{n},i\neq 0,n>c_{i},r,$ (34)
$c_{i}(y_{in}-y_{inr})\leq\sum_{m=1}^{n-1}w_{imr},$ $\forall i\in C_{n},i\neq 0,n,r,$ (35)
$\alpha_{inrl}\leq\lambda_{inl},$ $\forall i\in C_{n},i\neq 0,n,r,l,$ (36)
$\alpha_{inrl}\leq w_{inr},$ $\forall i\in C_{n},i\neq 0,n,r,l,$ (37)
$\sum_{l}\lambda_{inl}\leq 1,$ $\forall i\in C_{n},i\neq 0,n,$ (38)
$lp_{in}+\sum_{l}pm\lambda_{inl}l\leq mp_{in}$ $\forall i\in C_{n},i\neq 0,n$ (39)
Constraints 19, 20, 30 ensure that once an alternative is not in the choice set of a customer, it will never be chosen by the customer.
Constraints 33, 34, 35 ensure that it is possible to have $y_{inr}=1$ when all the customers before the $n_{i}$ have been served, and there exist enough quantity of alternative $i$ for the $n_{i}$ customer.
The above constraints do not offer any other available alternative to the customer. Constraints associated with discounted regret, 25,26, 27, 28, and 29 help find the available alternatives with the least anticipated regret in the choice set of a customer. Constraints 30 , 31, 32 address the issue of picking up the alternative with the least regret while it has enough quantity to be allocated. The rest of the constraints are the same as in the previous model. Finally, constraints 19, 20, 33, 34, 35, 36, 37, 39 were taken from pacheco2016new .
4 Computational experiments
In this section, we employ IBM ILOG CPLEX Optimization to execute the models and assess their performance in solving the numerical experiments.
In design of the experiments, we included three alternatives in the choice set for each customer. One of them represents the no-buy option, while the other two alternatives are left unlabeled. Regret arises due to the price attribute and unobserved factors. We did not consider the socio-economic characteristics of the customers in this experiment, i.e., $ER_{ijnr}=0,\ \forall i,j,n,r$. Specifically, price is the only attribute for both alternatives. Customers are homogeneous in their preferences, so $\beta_{np}=-1,\ \forall n.$ Also, $v_{onr}$ and $v_{nr}$ are positive random draws from $Gumbel(0,1)$ distribution. Price has six levels ranging from $1$ to $4.5$, with $pm=0.5$. The bounds are set to be arbitrarily large enough numbers. As for the RUM-based MILP modeled by pacheco2017integrating , the overall utility is solely associated with the price attribute and the unobserved utility, which is a random draw from $Gumbel(0,1)$.
Firstly, we solved both capacitated and uncapacitated RRM-based MILP models to optimality. We considered $r=4$ draws for each customer in both models. Table 1 summarizes the performance of the models. To check if capacity constraints are active, examine the instance with $11$ customers. There are five units of each of alternatives. Thus, there is only the opt-out choice for one of the customers, which shows itself in the gap between the uncapacitated and capacitated solution.
Table 1: Performance of Capacitated and Uncapacitated RRM-based MILP Models
No. Customer Capacity Time (s) Solution
Alt. 2 Alt. 3 Cap Uncap Cap Uncap
10 5 5 171 2.2 45 45
11 5 5 216 3.4 45 49.5
12 6 6 206 3.5 54 54
13 7 6 357 3.6 58.5 58.5
14 7 7 787 9.5 63 63
15 8 7 1320 5.28 67.5 67.5
Also, it is important to note that despite having the same number of customers and equal capacity for the alternatives (except in the case of 11 customers), the uncapacitated model was solved more quickly. This could be attributed to the fact that the capacitated version involves additional constraints to adhere to capacity restrictions, leading to delays in solving the model. Additionally, we conclude that the opt-out option was never chosen given the increasing revenue trend while the capacity of each of alternatives has increased by one unit.
Given that the customers and alternatives are identical, the number of feasible solutions and iterations tends to be high due to possible symmetries. The models involve a substantial number of variables and constants, even for a relatively small number of customers and draws, as depicted in Table 2. The combinatorial nature of these models necessitates the application of heuristic methods to address computational challenges.
Table 2: Structure and Performance of Capacitated and Uncapacitated RRM-based MILP Models
No. Customer Constraints Variables Iterations Nodes
Cap Uncap Cap Uncap Cap Uncap Cap Uncap
10 5412 2520 1910 1530 10291738 5471 175717 0
11 5991 2772 2112 1683 5466747 7312 67791 6355
12 6532 3024 2304 1836 8627555 8752 147674 0
13 7077 3276 2496 1989 6352511 8538 67486 0
14 7622 3528 2688 2142 12983867 343831 222918 0
15 8167 3780 2880 2295 18771065 13401 169710 0
Table 3 holds particular significance. In the RRM-based model, higher price levels have been selected, in contrast to the RUM-based model where lower price levels are favored. This discrepancy highlights the impact of different choice behavior theories on results, emphasizing the importance of utilizing these theories appropriately in demand and supply studies. Applying RUM-based models where customers are regret minimizers can lead to huge profit losses.
Table 3: Performance of Uncapacitated RRM and RUM-based MILP Models
No. Customers Time, RUM (s) Time, RRM (s) Solution, RUM Solution, RRM Gap(%)
10 0.08 2.2 7.125 45 84
11 0.02 3.4 21.125 49.5 57
12 0.08 3.5 10 54 81
13 0.06 3.6 8.25 58.5 85
14 0.06 9.5 7.25 63 88
15 0.08 5.28 15 67.5 77
The structure of the RRM-based model, which compares alternatives pairwise, introduces many symmetries in this particular experiment, thereby increasing the number of iterations, as shown in Table 4. Additionally, there are more variables and constraints in the RRM-based MILP compared to its RUM-based counterpart signaling potential application of heauristics and more sophisticated models for RRM-based MILPs to tackle large size problems.
Table 4: Structure and Performance of Uncapacitated RRM and RUM-based MILP Models
No. Customer Constraints Variables No. Iter
RUM-based RRM-based RUM-based RRM-based RUM-based RRM-based
10 1570 2520 1050 1530 26 5471
11 1727 2772 1155 1683 17 7312
12 1884 3024 1260 1836 23 8752
13 2041 3276 1365 1989 27 8538
14 2198 3528 1470 2142 12 343831
15 2355 3780 1575 2295 47 13401
In conclusion, this experiment demonstrated the feasibility of our models and underscored the significance of studying different consumer choice behaviors, as they can impact revenue, pricing, and capacity allocation policies. Our results demonstrate the superiority of our RRM-based MILP in revenue maximization when customers are regret minimizers.
5 Conclusion
In conclusion, cutting-edge Revenue Management (RM) techniques, adept at handling the intricate dynamics of supply and demand, leverage Discrete Choice Models (DCMs) as superior demand estimators rooted in consumer purchasing behavior. These models are seamlessly integrated into Mixed-Integer Linear Programming (MILP) frameworks, drawing inspiration from Random Utility Maximization (RUM) theory and the more recent Random Regret Minimization (RRM) theory. Notably, RUM-based DCMs have been the go-to choice in the RM landscape.
This study contributes by formulating MILP models tailored for both capacitated and uncapacitated supply scenarios, with a specific focus on customers as regret minimizers based on RRM theory. The results not only validate the viability of these models but also underscore the substantial impact of varied customer behavior interpretations on revenue, pricing, and customer choices.
Future research endeavors may entail tackling large-scale datasets to solve RUM-based and RRM-based MILP models, shedding light on their real-world applicability. In the face of potential computational challenges, especially with the heightened complexity of RRM-based MILP models, adopting Benders-type heuristics, column generation, or meta heuristics emerge as a pragmatic solution.
Furthermore, recognizing that both RUM-based and RRM-based MILPs yield extreme values for prices, utility, and regret, the development of bi-objective MILPs is proposed. Introducing a second objective, aimed at maximizing utility or minimizing regret, could offer deeper insights. Exploring Pareto optimal frontiers in such models promises nuanced information about customer satisfaction, empowering suppliers with a more informed approach to decision-making in Revenue Management.
6 Declaration of competing interest
The author affirms that there are no identifiable conflicting financial interests or personal associations that might have seemed to impact the work described in this article.
7 Data and Code availability
Data and code are available upon request.
8 Acknowledgments
We acknowledge Grenoble Alpes university IDEX scholarship. Dr. Pierre Lemaire and Dr. Iragael Joly provided important support.